OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a19 = -92, a20 = 6

OpenStudy (cwrw238):

common difference = 6 - (-92) nth term = a + (n-1)d now u need to find to find a can u do this?

OpenStudy (anonymous):

is a -1856?

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

so now you've git it plug in a = -1856 and d = 98 into the general formula

OpenStudy (anonymous):

so the formula would be -1856 + 98(n - 1)

OpenStudy (anonymous):

or n+1

OpenStudy (anonymous):

?

OpenStudy (cwrw238):

its n - 1

OpenStudy (anonymous):

is it always -1?

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

check out the a20 term -1856 +(20-1)98 should equal 6

OpenStudy (anonymous):

okay thanks. can you help me with a few more?

OpenStudy (cwrw238):

no sorry i gotta go right now

OpenStudy (anonymous):

okay thanks anyway